Question:

The workers' union at a particular university is quite strong. About 96% of all workers employed by the university belong to the workers' union. Recently, the workers went on strike, and now a local TV station plans to interview5 workers (chosen at random) at the university to get their opinions on the strike. What is the probability that exactly 3 of the workers interviewed are union members?

Round your response to at least three decimal places. (If necessary, consult a list of formulas.)
Answer #1

5-3. þ = 96% =0-96 ,ne 5 Binomial distribution PIX) = () px (1-P)*** P(x= 3) = 0) (0.963 (1–0.96) = 0·014155 ~ 10.0142 that e